OpenStudy (gabylovesyou):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (goformit100):



OpenStudy (gabylovesyou):

what do i do ??

OpenStudy (compassionate):

Ohai Gabby.

OpenStudy (gabylovesyou):

hai lol

OpenStudy (compassionate):

Looks like you need help with math

OpenStudy (gabylovesyou):

yup :P

OpenStudy (compassionate):

Well, what seems to be the issue?

OpenStudy (gabylovesyou):

i dont know what to do here ^^^

OpenStudy (compassionate):

Are you aware how to graph?

OpenStudy (gabylovesyou):

no ;/

OpenStudy (compassionate):

No problem. If I gave you the points (2, 3) could you graph it?

OpenStudy (gabylovesyou):

yes :)

OpenStudy (compassionate):

Okay, good. Are you aware what slope is?

OpenStudy (gabylovesyou):

yes :)

OpenStudy (compassionate):

Okay good. You know all that you need to do in order too solve this problem. All you have to do is apply it in a slightly different way.

OpenStudy (compassionate):

Are you aware how to use y=mx+b?

OpenStudy (gabylovesyou):

no ;?

OpenStudy (compassionate):

Okay, well y=mx+b is basically a quick way to find the equation when given x, y, and the slope.

jimthompson5910 (jim_thompson5910):

y = mx+b y = -6x+b ... Plug in the given slope -3 = -6(-2)+b ... Plug in the given point Now solve for b